# Efficient Methods for Non-stationary Online Learning

### v2

@@ C.4 Proof of Lemma 5
$\sum_{\tau=i}^{t}f_{\tau}(\mathbf{x}_{\tau})-\sum_{t=i}^{t}f_{\tau}(\mathbf{u})$ ${}\leq\sum_{\tau=i}^{t}g_{\tau}(\mathbf{x}_{\tau})-\sum_{t=i}^{t}g_{\tau}(\mathbf{u})\leq\sum_{\tau=i}^{t}\langle\nabla g_{\tau}(\mathbf{y}_{\tau}),\mathbf{y}_{\tau}-\mathbf{u}\rangle$
@@
The second inequality makes use of $\sqrt{a+b}\leq\sqrt{a}+\sqrt{b}$ and $\sqrt{ab}\leq(a^{2}+b^{2})/2$. The last inequality holds by $(a+b+c)^{2}\leq 3(a^{2}+b^{2}+c^{2})$. ∎
@@ C.5 Proof of Lemma 6
For interval $[s_{m},s_{m+1}-1]$, there must exist an interval $[i,j]\in\widetilde{\mathcal{C}}$ such that $[s_{m},s_{m+1}-1]\subseteq[i,j]$ with $i=s_{m}$. Therefore, we can apply Eq. (34) presented in Lemma 5 to upper bound the regret during $[s_{m},s_{m+1}-1]$ for any comparator $\mathbf{u}\in\mathcal{X}$,
@@ C.6 Proof of Lemma 7
We assume that the $m$-th base-learner is initialized at the beginning of interval $[i,j]\in\widetilde{\mathcal{C}}$, in other words, $i=s_{m}$, where $s_{m}$ denotes the $m$-th marker. Before time stamp $i$, the schedule has registered $m-1$ markers, i.e., from $s_{1}$ to $s_{m-1}$. By Lemma 6, we can lower bound the cumulative loss of any comparator $\mathbf{u}\in\mathcal{X}$ as,
$\sum_{\tau=s_{1}}^{s_{m}-1}f_{\tau}(\mathbf{u})\geq\frac{1}{4}\sum_{k=1}^{m-1}C_{a}\geq\frac{C_{1}}{4}(m-1).$
@@ D.1 Projection-Efficient Variant for the Method of Cutkosky (2020)
Unlike the meta-base type algorithm (such as AOA (Zhang et al., 2020)) which combines decisions using a weighted vector $\bm{p}_{t}\in\Delta_{N}$ from the meta-algorithm as $\mathbf{x}_{t}=\sum_{i=1}^{N}p_{t,i}\mathbf{x}_{t,i}$utkosky (2020) aggregates the decisions from base-learners sequentially, submitting the decision as $\mathbf{x}_{t}=\sum_{i=1}^{N}\mathbf{x}_{t,i}$ with the number of base-learners $N=\mathcal{O}(\log T)$. Their method relies in the specific beneficial properties of parameter-free online learners (Cutkosky and Orabona, 2018). Specifically, with $\mathbf{x}_{t}^{(k-1)}=\sum_{i=1}^{k-1}\mathbf{x}_{t,i}$ representing the aggregated decisions of the first $k-1$ base-learners, the $k$-th base-learner then makes a decision $(\mathbf{y}_{t,k},z_{t,k})$ on a $(d+1)$-dimension lifted domain $V_{t}^{(k)}$ defined as
@@
If we choose $\mathcal{Y}$ to be an Euclidean ball, the corresponding lifted domain $\widehat{V}_{t}^{(k)}$ becomes a cylinder, which simplifies the projection for the base-learners. In this approach, we only require the $N$-th base-learner to project her decision onto the lifted domain constructed over $\mathcal{X}$, to ensure the final submitted decision $\mathbf{x}_{t}=\mathbf{x}_{t}^{(N)}=\sum_{i=1}^{N}\mathbf{x}_{t,i}\in\mathcal{X}$ lying within $\mathcal{X}$. After submitting $\mathbf{x}_{t}$, we send the gradient $\mathbf{g}_{t}=\nabla f_{t}(\mathbf{x}_{t})$ to the method of Cutkosky (2020) with the aforementioned revision. Thus, our reduction mechanism can improve efficiency by requiring only $1$ projection onto the complex lifted domain associated with $\mathcal{X}$ (as defined in (61) for $k=N$), while performing $\mathcal{O}(\log T)$ projections onto simpler lifted domains (as defined in (62) for $k=1,\cdots,N-1$).
@@
Even with our reduction mechanism applied, the variant of  Cutkosky (2020)’s method still necessitates at least $1$ projection onto the complex lifted domain associated with $\mathcal{X}$, which is more complicated than the simple projection onto $\mathcal{X}$ required by our method. Furthermore, although performing $\mathcal{O}(\log T)$ projections onto simpler lifted domains is easy and feasible, it is still more complex than the simple rescaling operation our method uses in order to projections onto the (surrogate) Euclidean ball.
@@ D.2 Key Lemmas
Under Assumptions 12, and 3, setting the step size pool as =$\mathcal{H}=\big{\{}\eta_{j}=2^{j-1}\sqrt{5D^{2}/(1+8LGDT)}\mid j\in[N]\big{\}}$ with $N=\lceil 2^{-1}\log_{2}((5D^{2}+2D^{2}T)(1+8LGDT)/(5D^{2}))\rceil+1$ and $\varepsilon_{i,\tau}=\sqrt{(\ln N)/(1+D^{2}\sum_{s=i}^{\tau-1}\|\nabla g_{s}(\mathbf{y}_{s})\|_{2}^{2})}$, assume that the $m$-th base-learner described in Algorithm 3 is initialized at the beginning of an interval $[i,j]\in\widetilde{\mathcal{C}}$, then this base-learner ensures the following dynamic regret
@@
Combining above lemma with the analysis of the meta-algorithm Adapt-ML-Prod (see Lemma 4), we can obtain the interval dynamic regret for intervals in the problem-dependent geometric covering intervals $\widetilde{\mathcal{C}}$.
@@
Under Assumptions 12, and 3, for any interval $[s_{m},s_{m+1}-1]$ determined by two consecutive intervals $s_{m}$ and $s_{m+1}$, where we denote by $s_{m}$ the $m$-th marker, Algorithm 3 ensures that for any comparators $\mathbf{u}_{s_{m}},\dots,\mathbf{u}_{s_{m+1}-1}\in\mathcal{X}$, we have
@@ D.3 Proof of Theorem 5
The proof closely follows that of Theorem 4, with only the key steps outlined here. For a comprehensive understanding of the techniques, readers may check the proofs in Appendix C.2, where the problem setting is simpler with only a fixed comparator conisdered.
@@
For the interval $[r,s_{p}-1]$, we can directly bound the regret by the threshold of the cumulative loss and obtain that $\sum_{t=r}^{s}f_{t}(\mathbf{x}_{t})-\sum_{t=q}^{s}f_{t}(\mathbf{u}_{t})\leq\sum_{t=s_{p-1}}^{s_{p}-1}f_{t}(\mathbf{x}_{t})\leq C_{p-1}+GD$, due to the threshold mechanism and the non-negative property of loss functions. The threshold generating function specified in Eq. (19) further indicates $C_{p-1}\leq\mathcal{O}\left(\log(F_{T}+P_{T})\right)$.
@@ D.5 Proof of Lemma 9
$\sum_{\tau=i}^{t}f_{\tau}(\mathbf{x}_{\tau})-f_{\tau}(\mathbf{u}_{\tau})$ $\leq\sum_{\tau=i}^{t}\langle\nabla g_{\tau}(\mathbf{y}_{\tau}),\mathbf{y}_{\tau}-\mathbf{u}_{\tau}\rangle$
@@
$\sum_{\tau=i}^{t}\langle\nabla g_{t}(\mathbf{y}_{t}),\mathbf{y}_{t}-\mathbf{y}_{t,m}\rangle\leq 4D\left(3\sqrt{\ln(1+2m)}+\mu_{t}\right)\sqrt{LF_{[i,t]}^{\mathbf{x}}}+18GD\ln(1+2m)+6GD\mu_{t},$
@@ D.6 Proof of Lemma 10
$\sum_{t\in I}f_{\tau}(\mathbf{x}_{\tau})\leq 2\sum_{t\in I}f_{\tau}(\mathbf{u}_{\tau})+(126L+5)DP_{I}+\frac{1}{2}\mathcal{G}(k),$
@@
$2F^{\mathbf{u}}_{I}+(126L+5)DP_{I}\geq\sum_{t\in I}f_{t}(\mathbf{x}_{t})-\frac{1}{2}\mathcal{G}(k)\geq C_{k}-\frac{1}{2}\mathcal{G}(k)=\frac{1}{2}C_{k},$
@@ D.7 Proof of Lemma 11
We assume that the $m$-th base-learner is initialized at the beginning of interval $[i,j]\in\widetilde{\mathcal{C}}$ (in other words, $i=s_{m}$), where $s_{m}$ denotes the $m$-th marker. Before time stamp $i$, the schedule has registered $m-1$ markers, i.e., from $s_{1}$ to $s_{m-1}$. By Lemma 10, for any comparators, we have
@@ E.1.1 Assumptions
      The controller and transforming matrices and bounded, i.e., $\lVert K\rVert\leq\kappa$ and $\lVert H\rVert,\lVert H^{-1}\rVert\leq\kappa$.
@@ E.1.2 Proof of Theorem 6
The proof mainly follows the one of Scream.Control. We present the essential steps to demonstrate the application of efficient reduction here and refer the interested readers to Appendix C.2.3 in Zhao et al. (2022a) for comprehensive proof.
@@
where $\lambda=(H+2)^{2}L_{f}$ is a constant. The first inequality is by the coordiante-Lipschitz continuity of truncated function $f_{t}(\cdot)$. The third inequality is by the reduction mechanism and the final inequality is by $\lVert M_{t-1}-M_{t}\rVert_{\operatorname{F}}\leq\lVert M_{t-1}^{\prime}-M_{t}^{\prime}\rVert_{\operatorname{F}}$ derived from the nonexpanding property of projection (Nemirovski et al., 2009) and that one can verify in general this property holds for nearest point projection in Hilbert space.
Remind that Scream.Control employed in Algorithm 4 aim at minimizing the dynamic regret with switching-cost in domain $\mathcal{M}^{\prime}$, which can guarantee $\widetilde{\mathcal{O}}(\sqrt{T(1+P_{T})})$ regret bound by Theorem 4 in Zhao et al. (2022a). Thus, by taking into account that $\lVert\nabla g_{t}\left(M_{t}^{\prime}\right)\rVert_{\operatorname{F}}\leq\lVert\nabla\widetilde{f}_{t}\left(M_{t}\right)\rVert_{\operatorname{F}}$ by the efficient reduction mechanism, which is true under non-stochastic control setting since the algorithm optimizes the linearized loss and employs the Frobenius norm as the projection distance metric, we can derive our result. ∎
@@ E.2.1 Key Lemmas
$\sum_{t=r}^{s}\operatorname{tr}\left(\nabla g_{t}(\widehat{\mathbf{P}}^{s}_{t})\cdot\widehat{\mathbf{P}}^{s}_{t,i}\right)-\sum_{t=r}^{s}\operatorname{tr}\left(\nabla g_{t}(\widehat{\mathbf{P}}^{s}_{t})\cdot\mathbf{P}\right)\leq\mathcal{O}\left(\sqrt{k\cdot\lvert I\rvert}\right),$
The above claim can be verified by the proof in Appendix. B of Nie et al. (2016) together with $\lVert\nabla g_{t}(\widehat{\mathbf{P}}^{s}_{t})\rVert_{\operatorname{F}}\leq\lVert\nabla f_{t}(\widehat{\mathbf{P}}_{t})\rVert_{\operatorname{F}}$, which can be verified by noticing that the loss function $f_{t}(\mathbf{P})$ is coordiante-wise linear with $\mathbf{P}$ and we use Frobenius norm as the distance metric.
@@
Let $\mathbf{P}^{\prime}\in\mathbb{R}^{d\times d}$ be a symmetric matrix, with eigenvalues $\sigma_{1}^{\prime},\dots,\sigma_{d}^{\prime}$ and associated eigenvectors $\mathbf{v}_{1}^{\prime},\dots,\mathbf{v}_{d}^{\prime}$ Its projection $\mathbf{P}=\Pi_{\widehat{\mathcal{P}}_{k}}[\mathbf{P}^{\prime}]$ onto the domain $\widehat{\mathcal{P}}_{k}$ with respect to the Frobenius norm, is the unique feasible matrix which has the same eigenvectors as $\mathbf{P}^{\prime}$, with the associated eigenvalues $\sigma_{1},\dots,\sigma_{d}$ satisfying:

### v3

@@ C.4 Proof of Lemma 5
$\sum_{\tau=i}^{t}f_{\tau}(\mathbf{x}_{\tau})-\sum_{\tau=i}^{t}f_{\tau}(\mathbf{u})$ ${}\leq\sum_{\tau=i}^{t}g_{\tau}(\mathbf{x}_{\tau})-\sum_{t=i}^{t}g_{\tau}(\mathbf{u})\leq\sum_{\tau=i}^{t}\langle\nabla g_{\tau}(\mathbf{y}_{\tau}),\mathbf{y}_{\tau}-\mathbf{u}\rangle$
@@
The second inequality makes use of $\sqrt{a+b}\leq\sqrt{a}+\sqrt{b}$ and $\sqrt{ab}\leq(a+b)/2$. The last inequality holds by $(a+b+c)^{2}\leq 3(a^{2}+b^{2}+c^{2})$. ∎
@@ C.5 Proof of Lemma 6
For interval $s_{m},s_{m+1}-1]$, there must exist an interval $i,j]\in\widetilde{\mathcal{C}}$ such that $s_{m},s_{m+1}-1]\subseteq[i,j]$ with $i=s_{m}$. Therefore, we can apply Eq. (34) presented in Lemma 5 to upper bound the regret during $s_{m},s_{m+1}-1]$ for any comparator $\mathbf{u}\in\mathcal{X}$,
@@ C.6 Proof of Lemma 7
We assume that the $m$-th base-learner is initialized at the beginning of interval $i,j]\in\widetilde{\mathcal{C}}$, in other words, $i=s_{m}$, where $s_{m}$ denotes the $m$-th marker. Before time stamp $i$, the schedule has registered m1m-1 markers, i.e., from $s_{1}$ to $s_{m-1}$. By Lemma 6, we can lower bound the cumulative loss of any comparator $\mathbf{u}\in\mathcal{X}$ as,
$\sum_{\tau=s_{1}}^{s_{m}-1}f_{\tau}(\mathbf{u})\geq\frac{1}{4}\sum_{k=1}^{m-1}C_{k}\geq\frac{C_{1}}{4}(m-1).$
@@ D.1 Projection-Efficient Variant for the Method of Cutkosky (2020)
Unlike the meta-base type algorithm (such as AOA (Zhang et al., 2020)) which combines decisions using a weighted vector $\bm{p}_{t}\in\Delta_{N}$ from the meta-algorithm as $\mathbf{x}_{t}=\sum_{i=1}^{N}p_{t,i}\mathbf{x}_{t,i}$Cutkosky (2020) aggregates the decisions from base-learners sequentially, submitting the decision as $\mathbf{x}_{t}=\sum_{i=1}^{N}\mathbf{x}_{t,i}$ with the number of base-learners $N=\mathcal{O}(\log T)$. Their method relies on the specific beneficial properties of parameter-free online learners (Cutkosky and Orabona, 2018). Specifically, with $\mathbf{x}_{t}^{(k-1)}=\sum_{i=1}^{k-1}\mathbf{x}_{t,i}$ representing the aggregated decisions of the first k1k-1 base-learners, the $k$-th base-learner then makes a decision $(\mathbf{y}_{t,k},z_{t,k})$ on a $(d+1)$-dimensional lifted domain $V_{t}^{(k)}$ defined as
@@
If we choose $\mathcal{Y}$ to be an Euclidean ball, the corresponding lifted domain $\widehat{V}_{t}^{(k)}$ becomes a cylinder, which simplifies the projection for the base-learners. In this approach, we only require the $N$-th base-learner to project her decision onto the lifted domain constructed over $\mathcal{X}$, to ensure the final submitted decision $\mathbf{x}_{t}=\mathbf{x}_{t}^{(N)}=\sum_{i=1}^{N}\mathbf{x}_{t,i}\in\mathcal{X}$ to lie within $\mathcal{X}$. After submitting $\mathbf{x}_{t}$, we send the gradient $\mathbf{g}_{t}=\nabla f_{t}(\mathbf{x}_{t})$ to the method of Cutkosky (2020) with the aforementioned revision. Thus, our reduction mechanism can improve efficiency by requiring only $1$ projection onto the complex lifted domain associated with $\mathcal{X}$ (as defined in (61) for $k=N$), while performing $\mathcal{O}(\log T)$ projections onto simpler lifted domains (as defined in (62) for $k=1,\cdots,N-1$).
@@
Even with our reduction mechanism applied, the variant of  Cutkosky (2020)’s method still necessitates at least $1$ projection onto the complex lifted domain associated with $\mathcal{X}$, which is more complicated than the simple projection onto $\mathcal{X}$ required by our method. Furthermore, although performing $\mathcal{O}(\log T)$ projections onto simpler lifted domains is easy and feasible, it is still more complex than the simple rescaling operation our method uses for projections onto the (surrogate) Euclidean ball.
@@ D.2 Key Lemmas
Under Assumptions 12, and 3, setting the step size pool as $\mathcal{H}=\big{\{}\eta_{j}=2^{j-1}\sqrt{5D^{2}/(1+8LGDT)}\mid j\in[N]\big{\}}$ with $N=\lceil 2^{-1}\log_{2}((5D^{2}+2D^{2}T)(1+8LGDT)/(5D^{2}))\rceil+1$ and $\varepsilon_{i,\tau}=\sqrt{(\ln N)/(1+D^{2}\sum_{s=i}^{\tau-1}\|\nabla g_{s}(\mathbf{y}_{s})\|_{2}^{2})}$, assume that the $m$-th base-learner described in Algorithm 3 is initialized at the beginning of an interval $i,j]\in\widetilde{\mathcal{C}}$, then this base-learner ensures the following dynamic regret
@@
Combining the above lemma with the analysis of the meta-algorithm Adapt-ML-Prod (see Lemma 4), we can obtain the interval dynamic regret for intervals in the problem-dependent geometric covering intervals $\widetilde{\mathcal{C}}$.
@@
Under Assumptions 12, and 3, for any interval $s_{m},s_{m+1}-1]$ determined by two consecutive intervals $s_{m}$ and $s_{m+1}$, where we denote by $s_{m}$ the $m$-th marker, Algorithm 3 ensures that for any comparators $\mathbf{u}_{s_{m}},\dots,\mathbf{u}_{s_{m+1}-1}\in\mathcal{X}$, we have
@@ D.3 Proof of Theorem 5
The proof closely follows that of Theorem 4, with only the key steps outlined here. For a comprehensive understanding of the techniques, readers may check the proofs in Appendix C.2, where the problem setting is simpler with only a fixed comparator considered.
@@
For the interval $r,s_{p}-1]$, we can directly bound the regret by the threshold of the cumulative loss and obtain that $\sum_{t=r}^{s_{p}-1}f_{t}(\mathbf{x}_{t})-\sum_{t=r}^{s_{p}-1}f_{t}(\mathbf{u}_{t})\leq\sum_{t=s_{p-1}}^{s_{p}-1}f_{t}(\mathbf{x}_{t})\leq C_{p-1}+GD$, due to the threshold mechanism and the non-negative property of loss functions. The threshold generating function specified in Eq. (19) further indicates $C_{p-1}\leq\mathcal{O}\left(\log(F_{T}+P_{T})\right)$.
@@ D.5 Proof of Lemma 9
$\sum_{\tau=i}^{t}f_{\tau}(\mathbf{x}_{\tau})-\sum_{\tau=i}^{t}f_{\tau}(\mathbf{u}_{\tau})$ $\leq\sum_{\tau=i}^{t}\langle\nabla g_{\tau}(\mathbf{y}_{\tau}),\mathbf{y}_{\tau}-\mathbf{u}_{\tau}\rangle$
@@
$\sum_{\tau=i}^{t}\langle\nabla g_{\tau}(\mathbf{y}_{\tau}),\mathbf{y}_{\tau}-\mathbf{y}_{\tau,m}\rangle\leq 4D\left(3\sqrt{\ln(1+2m)}+\mu_{t}\right)\sqrt{LF_{[i,t]}^{\mathbf{x}}}+18GD\ln(1+2m)+6GD\mu_{t},$
@@ D.6 Proof of Lemma 10
$\sum_{t\in I}f_{t}(\mathbf{x}_{t})\leq 2\sum_{t\in I}f_{t}(\mathbf{u}_{t})+(126L+5)DP_{I}+\frac{1}{2}\mathcal{G}(m),$
@@
$2F^{\mathbf{u}}_{I}+(126L+5)DP_{I}\geq\sum_{t\in I}f_{t}(\mathbf{x}_{t})-\frac{1}{2}\mathcal{G}(m)\geq C_{m}-\frac{1}{2}\mathcal{G}(m)=\frac{1}{2}C_{m},$
@@ D.7 Proof of Lemma 11
We assume that the $m$-th base-learner is initialized at the beginning of interval $i,j]\in\widetilde{\mathcal{C}}$ (in other words, $i=s_{m}$), where $s_{m}$ denotes the $m$-th marker. Before time stamp $i$, the schedule has registered m1m-1 markers, i.e., from $s_{1}$ to $s_{m-1}$. By Lemma 10, for any comparators, we have
@@ E.1.1 Assumptions
      The controller and transforming matrices are bounded, i.e., $\lVert K\rVert\leq\kappa$ and $\lVert H\rVert,\lVert H^{-1}\rVert\leq\kappa$.
@@ E.1.2 Proof of Theorem 6
The proof mainly follows the one of Scream.Control. We present the essential steps to demonstrate the application of efficient reduction here and refer interested readers to Appendix C.2.3 in Zhao et al. (2022b) for comprehensive proof.
@@
where $\lambda=(H+2)^{2}L_{f}$ is a constant. The first inequality is by the coordinate-Lipschitz continuity of truncated function $f_{t}(\cdot)$. The third inequality is by the reduction mechanism and the final inequality is by $\lVert M_{t-1}-M_{t}\rVert_{\operatorname{F}}\leq\lVert M_{t-1}^{\prime}-M_{t}^{\prime}\rVert_{\operatorname{F}}$ derived from the nonexpanding property of projection (Nemirovski et al., 2009) and that one can verify in general this property holds for nearest point projection in Hilbert space.
Remind that Scream.Control employed in Algorithm 4 aims at minimizing the dynamic regret with switching-cost in domain $\mathcal{M}^{\prime}$, which can guarantee $\widetilde{\mathcal{O}}(\sqrt{T(1+P_{T})})$ regret bound by Theorem 4 in Zhao et al. (2022b). Thus, by taking into account that $\lVert\nabla g_{t}\left(M_{t}^{\prime}\right)\rVert_{\operatorname{F}}\leq\lVert\nabla\widetilde{f}_{t}\left(M_{t}\right)\rVert_{\operatorname{F}}$ by the efficient reduction mechanism, which is true under non-stochastic control setting since the algorithm optimizes the linearized loss and employs the Frobenius norm as the projection distance metric, we can derive our result. ∎
@@ E.2.1 Key Lemmas
[…]
The above claim can be verified by the proof in Appendix. B of Nie et al. (2016) together with $\lVert\nabla g_{t}(\widehat{\mathbf{P}}^{s}_{t})\rVert_{\operatorname{F}}\leq\lVert\nabla f_{t}(\widehat{\mathbf{P}}_{t})\rVert_{\operatorname{F}}$, which can be verified by noticing that the loss function $f_{t}(\mathbf{P})$ is coordinate-wise linear with $\mathbf{P}$ and we use Frobenius norm as the distance metric.
@@
Let $\mathbf{P}^{\prime}\in\mathbb{R}^{d\times d}$ be a symmetric matrix, with eigenvalues $\sigma_{1}^{\prime},\dots,\sigma_{d}^{\prime}$ and associated eigenvectors $\mathbf{v}_{1}^{\prime},\dots,\mathbf{v}_{d}^{\prime}$ . Its projection $\mathbf{P}=\Pi_{\widehat{\mathcal{P}}_{k}}[\mathbf{P}^{\prime}]$ onto the domain $\widehat{\mathcal{P}}_{k}$ with respect to the Frobenius norm, is the unique feasible matrix which has the same eigenvectors as $\mathbf{P}^{\prime}$, with the associated eigenvalues $\sigma_{1},\dots,\sigma_{d}$ satisfying: